OpenStudy (alex6799):

Short Answer: Write 32x^5y^5 with only one exponent. Use parentheses.

OpenStudy (alex6799):

@CGGURUMANJUNATH

OpenStudy (alex6799):

@LieutenantGeneral

OpenStudy (lieutenantgeneral):

32x^5y^5 = 0 or 800 x = 0 y = 0 so 32^5 = 32 * 5